# arXiv:2602.23464v1[cs.CR] 26 Feb 2026

### v1

@@ Abstract
Multi-scalar multiplication (MSM), $\mathsf{MSM}(\vec{P},\vec{x})=\sum_{i=1}^{n}x_{i}P_{i}$, is a dominant computational kernel in discrete-logarithm–based cryptography and often becomes a bottleneck for verifiers and other resource-constrained clients. We present 2G2T, a simple protocol for verifiably outsourcing MSM to an untrusted server. After a one-time keyed setup for fixed bases $\vec{P}$ (producing a public merged-bases vector $\vec{T}$ and client secret state), the server returns only two group elements per query: the claimed MSM result $A=\mathsf{MSM}(\vec{P},\vec{x})$ and an auxiliary value $B=\mathsf{MSM}(\vec{T},\vec{x})$. Verification requires a single length-$n$ field inner product and a constant number of group operations (two scalar multiplications and one addition), while the server performs two MSMs. In our Ristretto255 implementation, verification is up to $\sim 300\times$ faster than computing the MSM locally using a highly optimized MSM routine for $n$ up to $2^{18}$, and the server-to-client response is constant-size (two compressed group elements, 64 bytes on Ristretto255). Despite its simplicity and efficiency, 2G2T achieves statistical soundness: for any (even unbounded) adversarial server, the probability of accepting an incorrect result is at most $1/q$ per query, and at most $e/q$ over $e$ adaptive executions, in a prime-order group of size $q$.
@@ 1 Introduction
Challenges. Prior work on verifiable MSM outsourcing highlights two obstacles: (i) achieving a meaningful speedup while keeping statistical soundness extremely high, and (ii) avoiding lightweight checks that enforce only a consistency relation between returned values but do not bind the server to all $n$ terms of the instance. For example, Pinkas [18] gives a non-interactive, publicly verifiable statistical check that can yield a modest constant-factor speedup when one targets moderate soundness (on the order of tens of bits), but doing so typically increases the server’s response size to hundreds of group elements on 255-bit groups, and the speedup diminishes as the soundness target approaches “full-group” levels. At the other end of the spectrum, cheaper consistency-only strategies can fail outright against a malicious server: a server may compute an MSM over only a strict subset of indices (e.g., return $A=x_{1}P_{1}$) and still satisfy the verifier’s consistency relation, even though $A\neq\mathsf{MSM}(\vec{P},\vec{x})$ (see, e.g., [21]). Together, these examples underscore that verifiable MSM outsourcing is nontrivial: one must simultaneously achieve negligible soundness error and bind the server to the full MSM instance without reintroducing MSM-like verifier work.
@@
Our contribution: 2G2T. We present 2G2T, a statistically sound protocol for verifiable MSM outsourcing with constant-size server-to-client communication and fast client-side verification. [^1] The mnemonic 2G2T reflects (i) “two group elements to transfer”: the server returns two group elements per query; and (ii) it can seem “too good to be true” at first glance, as it combines constant-size communication, $O(1)$ verifier group operations, and $1/q$-level statistical soundness. After a one-time setup for fixed bases $\vec{P}$, the server answers each client’s query $\vec{x}$ with only two group elements: $A$ claimed to equal $\mathsf{MSM}(\vec{P},\vec{x})$ and an auxiliary value $B=\mathsf{MSM}(\vec{T},\vec{x})$ for a precomputed merged-bases vector $\vec{T}$. The client verifies using a single length-$n$ field inner product $s=\langle\vec{x},\vec{\rho}\rangle$ and a constant number of group operations (two scalar multiplications and one addition), while the server performs exactly two MSMs (only one more than computing $A$ alone). We prove that 2G2T is perfectly complete and statistically sound: for any (even computationally unbounded) adversarial server, the probability that the verifier accepts an incorrect $A$ is at most $1/q$ per execution, and at most $e/q$ over $e$ adaptive executions (Theorems 4.14.3).
Our Rust implementation over Ristretto255 confirms that 2G2T yields large practical gains in the large-$n$ regime: verification is up to $\sim 300\times$ faster than computing a highly optimized MSM, and up to $\sim 3{,}000\times$ faster than a naïve “$n$ scalar multiplications” baseline, for dimensions up to $n=2^{18}=262{,}144$ (with exact constants depending on hardware and implementation details). Finally, server-to-client communication is constant and minimal: the response consists of two compressed group elements $(A,B)$, i.e., 64 bytes in Ristretto255, independent of $n$. These efficiency gains come while providing “full-group” statistical integrity: the per-query soundness error $1/q$ corresponds to about $\log_{2}(q)$ bits of soundness (roughly $252$ bits for Ristretto255).
@@ 2 Threat model and security goals
  • Completeness. If $\mathcal{S}$ follows the protocol and returns correct values, then $\mathcal{C}$ accepts and outputs $Y=\mathsf{MSM}(\vec{P},\vec{x})$ for every query $\vec{x}$.
@@ 3 The 2G2T Protocol
Lemma 2(Uniqueness of an accepting scalar for an incorrect $A$)
@@
where the probability is over the client’s randomness in the precomputation of Definition 1.
@@
Thus, online server work is essentially one additional MSM relative to computing $A$ alone. In practice, implementations based on windowing (e.g., Pippenger-style bucket methods) can reuse the scalar decomposition of $\vec{x}$ across both MSMs, since the scalars are identical and only the bases differ. Consequently, the incremental cost of the second MSM is typically less than a full second pass of all scalar preprocessing.
@@
  • two scalar multiplications in $\mathbb{G}$ (compute $rA$ and $sQ$),
@@
The precomputation phase computes $\vec{T}=(T_{1},\dots,T_{n})$ with $T_{i}=rP_{i}+\rho_{i}Q$ for all $i$. This incurs $n$ scalar multiplications by a fixed scalar $r$ on the bases $\vec{P}$, $n$ scalar multiplications on the fixed base $Q$, and $n$ group additions. Because this phase is executed once and amortized over many outsourced queries, its cost is typically dominated by the online savings when many MSMs are outsourced over the same bases $\vec{P}$. Moreover, fixed-base techniques can accelerate the $Q$-multiplications, since $Q$ is constant.
@@ 6 Related work
The closest related line of work that explicitly targets verifiable multi-exponentiation/MSM in transparent groups is due to Pinkas [18], which gives a non-interactive, publicly verifiable statistical check with soundness parameter $2^{-\lambda}$. At a high level, Pinkas’ protocol attains public verifiability by returning a proof consisting of on the order of $\log q$ group elements and by performing verifier work that depends on the target soundness parameter $\lambda$. In contrast, 2G2T is a designated-verifier protocol with a one-time keyed setup whose per-query verification is dominated by a single length-$n$ field inner product plus $O(1)$ group operations, while achieving constant-size server-to-client communication (two group elements) and per-query soundness error $1/q$ in prime-order groups. We view these as complementary points in the design space: public verifiability with larger proofs versus designated-verifier verification with constant-size responses and “full-group” statistical soundness.
A second line of work considers MSM outsourcing for zero-knowledge proof generation. In particular, Xiong et al. [21] propose an “immediate verification” step based on a consistency relation between multiple values returned by the server. However, as stated, such a consistency check does not by itself force the server to incorporate all $n$ terms of the MSM instance: a malicious server can satisfy the stated relation while computing only a strict subset of indices (e.g., returning a partial sum such as $A=x_{1}P_{1}$), even though $A\neq\mathsf{MSM}(\vec{P},\vec{x})$ in general. Thus, in the standard adversarial-server model for outsourcing, the described immediate-check mechanism does not, on its own, imply soundness.
@@ 7 Performance evaluation
We implement a reference benchmark in Rust using the curve25519-dalek library over the Ristretto255 group. Ristretto255 is a natural instantiation for our target setting: it provides a prime-order group abstraction, it is transparent (no pairings, no SRS), and it is commonly used in Bulletproofs-style deployments. Throughout, we fix the public point $Q$ to the standard Ristretto basepoint so that the verifier’s fixed-base multiplication $sQ$ can leverage the library’s precomputed basepoint table. [^2] The group-only part of verification is already negligible in our measurements; choosing $Q$ as the basepoint simply reflects a realistic engineering choice for an always-fixed public $Q$.
@@
followed by a constant amount of group work ($rA$, sQsQ, one addition, one comparison). To reduce modular-reduction overhead in the scalar field, we implement a lazy-reduction inner product: we accumulate products in a wide (512-bit) accumulator and reduce modulo $q$ only once per block of terms. In the experiments of Figure 1, we use a block size of $128$, which keeps the wide accumulator safely within 512 bits while substantially reducing the number of reductions. This optimization directly targets the verifier bottleneck and is faithful to how a high-performance 2G2T verifier would be engineered.
@@
For context, Pinkas [18] proposes a publicly verifiable statistical check for MSMs in prime-order groups. In his analysis, when scalars have bit-length $m$ (typically $m\approx\log_{2}q$), verification reduces to computing an $n$-wide MSM with shorter $(\lambda+\log_{2}m)$-bit exponents, yielding an expected speedup of roughly $\frac{m}{\lambda+\log_{2}m}$. Over Ristretto255, one may take $m\approx\log_{2}q\approx 252$ and $\log_{2}m\approx 8$, which predicts a speedup of about $252/(64+8)\approx 3.5\times$ at $\lambda=64$. In contrast, matching 2G2T’s per-query soundness error $1/q$ corresponds to $\lambda\approx\log_{2}q$ (about $252$), in which case the predicted speedup essentially vanishes. Pinkas’ protocol also requires the server to return $m$ intermediate group elements (about 252252256256 for Ristretto255), whereas 2G2T returns only two group elements $(A,B)$. These protocols therefore occupy different points in the design space: Pinkas provides public verifiability with $\Theta(\log q)$ communication, while 2G2T achieves constant-size responses and very fast designated-verifier checks after a one-time keyed setup.
@@ 8 Conclusion
We introduced 2G2T, a lightweight protocol for verifiably outsourcing generic-base multi-scalar multiplication $\mathsf{MSM}(\vec{P},\vec{x})$ in transparent prime-order groups. Our analysis shows that 2G2T is perfectly complete and achieves full-group statistical soundness: even against an unbounded adversarial server, the probability of accepting an incorrect $A$ is at most $1/q$ per execution, and at most $e/q$ over $e$ adaptive executions, requiring only that $\mathbb{G}$ has prime order $q$. Empirically, a Rust implementation over Ristretto255 confirms that verification is extremely fast in the large-$n$ regime: for $n$ up to $2^{18}$, verification is up to $\sim 300\times$ faster than locally computing an optimized MSM, while the server-to-client response remains constant-size (two compressed group elements, 64 bytes on Ristretto255). Overall, 2G2T isolates a simple and practical primitive for integrity-preserving MSM acceleration with minimal communication and verifier overhead.

### v2

# 2G2T: Constant-Size, Statistically Sound MSM Outsourcing

@@ Abstract
Multi-scalar multiplication (MSM), $\mathsf{MSM}(\vec{P},\vec{x})=\sum_{i=1}^{n}x_{i}P_{i}$, is a dominant computational kernel in discrete-logarithm–based cryptography and often becomes a bottleneck for verifiers and other resource-constrained clients. We present 2G2T, a simple protocol for verifiably outsourcing MSM to an untrusted server. 2G2T is efficient for both parties: the server performs only two MSM computations and returns only two group elements to the client, namely the claimed result $A=\mathsf{MSM}(\vec{P},\vec{x})$ and an auxiliary group element $B$. Client-side verification consists of a single length-$n$ field inner product and only three group operations (two scalar multiplications and one group addition). In our Ristretto255 implementation, verification is up to $\sim 300\times$ faster than computing the MSM locally using a highly optimized MSM routine (for $n$ up to $2^{18}$). Moreover, 2G2T enables latency-hiding verification: nearly all verifier work can be performed while waiting for the server’s response, so once $(A,B)$ arrives the verifier completes the check with only one scalar multiplication and one group addition (both independent of $n$). Finally, despite its simplicity and efficiency, we prove that 2G2T achieves statistical soundness: for any (even unbounded) adversarial server, the probability of accepting an incorrect result is at most $1/q$ per query, and at most $e/q$ over $e$ adaptive executions, in a prime-order group of size $q$.
@@ 1 Introduction
Our proposed method: 2G2T. We present 2G2T, a statistically sound protocol for verifiable MSM outsourcing with constant-size server-to-client communication and fast client-side verification. [^1] The mnemonic 2G2T reflects (i) “two group elements to transfer”: the server returns two group elements per query; and (ii) that the protocol can seem “too good to be true” at first glance, as it combines constant-size communication, $O(1)$ verifier group operations, and $1/q$-level statistical soundness. After a one-time setup for fixed bases $\vec{P}$, the server answers a query $\vec{x}$ with only two group elements: AA claimed to equal $\mathsf{MSM}(\vec{P},\vec{x})$ and an auxiliary value $B=\mathsf{MSM}(\vec{T},\vec{x})$ for a precomputed public merged-bases vector $\vec{T}$. The client verifies using a single length-$n$ field inner product and $O(1)$ group operations (two scalar multiplications and one addition), while the server performs exactly two MSMs (only one more than computing AA alone).
Our Rust implementation over Ristretto255 confirms that 2G2T yields large practical gains in the large-$n$ regime: verification is up to $\sim 300\times$ faster than computing a highly optimized MSM, and up to $\sim 3{,}000\times$ faster than a naïve “$n$ scalar multiplications” baseline, for dimensions up to $n=2^{18}$. Moreover, verification is latency-hiding: essentially all verifier work (the inner product and one fixed-base scalar multiplication) can be performed while waiting for the server’s response, so once $(A,B)$ arrives the client performs only one scalar multiplication and one group addition to complete the check. We prove that 2G2T is perfectly complete and statistically sound: for any (even computationally unbounded) adversarial server, the probability that the verifier accepts an incorrect AA is at most $1/q$ per execution (roughly 252252 bits for Ristretto255), and at most $e/q$ over $e$ adaptive executions (Theorems 4.14.3).
Encrypted MSM (EMSM) and composability. The closest prior work that directly targets MSM outsourcing against an untrusted server while also hiding the scalar vector $\vec{x}$ is the recent EMSM line of work [1], motivated primarily by prover-side outsourcing where $\vec{x}$ is sensitive and must remain hidden from the server. In contrast, we focus on the common verifier-side regime where $\vec{x}$ is public (or already known to the server) and the bottleneck is simply evaluating large MSMs. This includes, for example, SNARK/Bulletproofs-style verification, where the MSM scalars are derived from the public transcript and public inputs, and deployments where the prover already holds $\vec{x}$ and can act as the server. In this public-scalar setting, 2G2T provides an extremely lightweight integrity layer: the client verifies using a single length-$n$ inner product plus $O(1)$ curve operations, and verification is naturally latency-hiding. [^2] As reported in [1], EMSM’s client time for large instances is dominated by a pre-query encryption step that must complete before the request can be sent, which leaves substantially less opportunity for latency-hiding.
When scalar privacy is required, 2G2T composes cleanly with EMSM: EMSM can provide coefficient hiding, while 2G2T supplies the integrity check. This avoids the additional EMSM execution used to harden against an untrusted server, reducing client-side work by up to a factor of two, without increasing the server’s MSM work or the number of group elements returned (relative to EMSM’s untrusted server setting). At the same time, the integrity guarantee is upgraded to statistical soundness, rather than relying on EMSM’s dual-LPN-style assumptions.
@@ 2 Threat model and security goals
  • Completeness. If $\mathcal{S}$ follows the protocol and returns correct values, then $\mathcal{C}$ accepts and outputs $A=\mathsf{MSM}(\vec{P},\vec{x})$ for every query $\vec{x}$.
@@ 3 The 2G2T Protocol
Lemma 2(Uniqueness of an accepting scalar for an incorrect AA)
@@
where the probability is over the client’s randomness in the setup and any internal randomness used by the (possibly randomized) server.
@@
Thus, online server work is essentially one additional MSM relative to computing AA alone. In practice, implementations based on windowing (e.g., Pippenger-style bucket methods) can reuse the scalar decomposition of $\vec{x}$ across both MSMs, since the scalars are identical and only the bases differ. Consequently, the incremental cost of the second MSM is typically less than a full second pass of all scalar preprocessing.
@@
  • two scalar multiplications in $\mathbb{G}$ (compute rArA and sQsQ),
@@
The precomputation phase computes $\vec{T}=(T_{1},\dots,T_{n})$ with $T_{i}=rP_{i}+\rho_{i}Q$ for all ii. This incurs $n$ scalar multiplications by a fixed scalar $r$ on the bases $\vec{P}$, $n$ scalar multiplications on the fixed base $Q$, and $n$ group additions. Because this phase is executed once and amortized over many outsourced queries, its cost is typically dominated by the online savings when many MSMs are outsourced over the same bases $\vec{P}$. Moreover, fixed-base techniques can accelerate the $Q$-multiplications, since $Q$ is constant.
@@ 6 Related work
The closest related line of work that explicitly targets verifiable multi-exponentiation/MSM in transparent groups is due to Pinkas [19], which gives a non-interactive, publicly verifiable statistical check with soundness parameter $2^{-\lambda}$. At a high level, Pinkas’ protocol attains public verifiability by returning a proof consisting of $O(\log q)$ group elements and by performing verifier work that depends on the target soundness parameter $\lambda$. In contrast, 2G2T is a designated-verifier protocol with a one-time keyed setup whose per-query verification is dominated by a single length-$n$ field inner product plus $O(1)$ group operations, while achieving constant-size server-to-client communication (two group elements) and per-query soundness error $1/q$ in prime-order groups. We view these as complementary points in the design space: public verifiability with larger proofs versus designated-verifier verification with constant-size responses and “full-group” statistical soundness.
A second line of work considers MSM outsourcing for zero-knowledge proof generation. In particular, Xiong et al. [22] propose an “immediate verification” step based on a consistency relation between multiple values returned by the server. However, such a consistency check does not by itself force the server to incorporate all $n$ terms of the MSM instance: a malicious server can satisfy the stated relation while computing only a strict subset of indices (e.g., returning a partial sum such as $A=x_{1}P_{1}$), even though $A\neq\mathsf{MSM}(\vec{P},\vec{x})$ in general. Thus, in the standard adversarial-server model for outsourcing, the described immediate-check mechanism does not, on its own, imply soundness.
@@ 7 Performance evaluation
We implement a reference benchmark in Rust using the curve25519-dalek library over the Ristretto255 group. Ristretto255 is a natural instantiation for our target setting: it provides a prime-order group abstraction, it is transparent (no pairings, no SRS), and it is commonly used in Bulletproofs-style deployments. Throughout, we fix the public point $Q$ to the standard Ristretto basepoint so that the verifier’s fixed-base multiplication sQsQ can leverage the library’s precomputed basepoint table. [^3] The group-only part of verification is already negligible in our measurements; choosing $Q$ as the basepoint simply reflects a realistic engineering choice for an always-fixed public $Q$.
@@
followed by a constant amount of group work (rArA, sQsQ, one addition, one comparison). To reduce modular-reduction overhead in the scalar field, we implement a lazy-reduction inner product: we accumulate products in a wide (512-bit) accumulator and reduce modulo $q$ only once per block of terms. In the experiments of Figure 1, we use a block size of 128128, which keeps the wide accumulator safely within 512 bits while substantially reducing the number of reductions. This optimization directly targets the verifier bottleneck and is faithful to how a high-performance 2G2T verifier would be engineered.
@@
For context, Pinkas [19] proposes a publicly verifiable statistical check for MSMs in prime-order groups. In his analysis, when scalars have bit-length $m$ (typically $m\approx\log_{2}q$), verification reduces to computing an $n$-wide MSM with shorter $(\lambda+\log_{2}m)$-bit exponents, yielding an expected speedup of roughly $\frac{m}{\lambda+\log_{2}m}$. Over Ristretto255, one may take $m\approx\log_{2}q\approx 252$ and $\log_{2}m\approx 8$, which predicts a speedup of about $252/(64+8)\approx 3.5\times$ at $\lambda=64$. In contrast, matching 2G2T’s per-query soundness error $1/q$ corresponds to $\lambda\approx\log_{2}q$ (about 252252), in which case the predicted speedup essentially vanishes. Pinkas’ protocol also requires the server to return $m$ intermediate group elements (about 252252256256 for Ristretto255), whereas 2G2T returns only two group elements $(A,B)$. These protocols therefore occupy different points in the design space: Pinkas provides public verifiability with $\Theta(\log q)$ communication, while 2G2T achieves constant-size responses and very fast designated-verifier checks after a one-time keyed setup.
@@ 8 Conclusion
We introduced 2G2T, a lightweight protocol for verifiably outsourcing generic-base multi-scalar multiplication $\mathsf{MSM}(\vec{P},\vec{x})$ in transparent prime-order groups. Our analysis shows that 2G2T is perfectly complete and achieves full-group statistical soundness: even against an unbounded adversarial server, the probability of accepting an incorrect AA is at most $1/q$ per execution, and at most $e/q$ over $e$ adaptive executions, requiring only that $\mathbb{G}$ has prime order $q$. Empirically, a Rust implementation over Ristretto255 confirms that verification is extremely fast in the large-$n$ regime: for $n$ up to $2^{18}$, verification is up to $\sim 300\times$ faster than locally computing an optimized MSM, while the server-to-client response remains constant-size (two compressed group elements, 64 bytes on Ristretto255). Overall, 2G2T isolates a simple and practical primitive for integrity-preserving MSM acceleration with minimal communication and verifier overhead.